Find the equation in standard form of the line perpendicular to 2x-3y=-5 and passing through (3, -2). Write in standard form with all integer coefficients.
Answer by RAY100(1637)

2x-3y=-5
put into form y=mx+b
subt 2x each side
-3y=-2x-5
divide both sides by (-3)
y=(2/3)x+5
m1 = 2/3, to be perpendicular m2= - 1/m1 = -3/2
writing the per ln eq in form y=mx+b
y= (-3/2)x +b
subst (3,-2)
(-2)= (-3/2) (3) =b
-2 = -9/2 +b
add 9/2 to both sides
-2 +9/2 =b
5/2=b
or y=mx+b
y= (-3/2)x +5/2
mult by 2
2y=-3x+5
add 3x to each side
3x+2y =5 ANSWER in standard form (Ax +By =C)
check 3(3)+2(-2) = 5 ok